the three angles of a triangle are (x+20),(x), and (x-50)degrees. Find the measure of each angle.
Answer by checkley77(12844)

(x+20),(x), and (x-50)
x+20+x+x-50=180
3x=180-20+50
3x=210
x=210/3
x=70 ans.
70+20=90 ans.
70-50=20 ans
Proof:
70+90+20=180
180=180